OpenStudy (anonymous):

solve the permutation of 45 objects chosen 6 at a time a 270 b720 c815,060 d 5,864,443,200 e 8,303,765,625

OpenStudy (anonymous):

i think itz 270 dk if im right

OpenStudy (amistre64):

what does that mean, chosen 6 at a time?

OpenStudy (amistre64):

45 pick 6 maybe?

OpenStudy (anonymous):

idk it was in the ?

OpenStudy (amistre64):

58644.... if i see it right

OpenStudy (anonymous):

45!/39!=45x44x43x42x41x40

OpenStudy (amistre64):

45.44.43.42.41.40

OpenStudy (amistre64):

6 doesnt divde 45; so picking 6 at a time sounds weird

OpenStudy (anonymous):

How many different ways can you permute 6 objects from a group of 45. \[45P6 = {45! \over (45-6)!} = {45! \over 39!} = 45\times 44 \times 43 \times 42 \times 41 \times 40 = \]